OpenStudy (anonymous):

What is the difference in volume between a 13 ft. cube and a 10 ft. cube? A. 1500 cubic ft. B. 1197 cubic ft. C. 3 cubic ft. D. 1300 cubic ft.

OpenStudy (zale101):

x^3= x cubed \((13ft)^3=13 ~ft.~cubed\) \((10ft)^3=10~ ft. ~cube\) Difference means subtract between those two volumes

OpenStudy (anonymous):

C?